The brand manager for a brand of toothpaste must plan a campaign designed to increase brand recognition. He wants to first determine the percentage of adults who have heard of the brand. How many adults must he survey in order to be 95​% confident that his estimate is within six percentage points of the true population​ percentage? Complete parts​ (a) through​ (c) below. ​a) Assume that nothing is known about the percentage of adults who have heard of the brand.

n=

statistics
Answer #1

Solution:

Given:

c = confidence level = 95% = 0.95

E = Margin of Error = within six percentage points of the true population ​percentage

E = 6%  

E = 0.06

p is unknown

thus p = 0.5

We have to find sample size n.

We need to find zc value for c=95% confidence level.

Find Area = ( 1 + c ) / 2 = ( 1 + 0.95) /2 = 1.95 / 2 = 0.9750

Look in z table for Area = 0.9750 or its closest area and find z value.

Area = 0.9750 corresponds to 1.9 and 0.06 , thus z critical value = 1.96

That is : Zc = 1.96

( Sample size is always rounded up)